# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v5

@@ Abstract
Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with various strategies of instruction augmentation, ultimately boosting model performance within specific domains. To date, the factors affecting FedDIT remain unclear, and existing instruction augmentation methods primarily focus on the centralized setting without considering distributed environments. Our experiments reveal that the cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT. In response, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based augmentation. At its core, the greedy selection procedure iteratively picks client centers that maximize the diversity and coverage of the instruction space while avoiding redundancy with previously selected centers. This ensures broad yet efficient coverage of the domain distribution across clients. For client-side computational efficiency and system scalability, FedDCA, the variant of FedDCA, utilizes heterogeneous encoders with server-side feature alignment. Extensive experiments across code, medical, financial, and mathematical domains substantiate the effectiveness of both methods, as well as plug-and-play capability. We further analyze privacy preservation against memory extraction attacks, showing that while privacy leakage risk is independent of augmented public data ratio, it decreases or converges as training progresses.
@@ 1 Introduction
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data McMahan et al. (2017); Ye et al. (2024b); Zhang et al. (2023b). This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization. Despite the establishment of various FedIT frameworks Ye et al. (2024b); Kuang et al. (2023); Zhang et al. (2023b), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications. For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts Zhang et al. (2024b).
Figure 1: Visualization of different instruction augmentation strategies in the mathematical domain. The Random Sampling strategy randomly samples instructions from the public dataset. Direct Retrieval performs dense retrieval from the public dataset based on the client’s local data distribution without considering the cross-client domain coverage. FedDCA strategically selects client centers to maximize cross-client domain coverage and then performs domain data retrieval. The t-SNE visualization shows FedDCA has a better domain coverage than other sampling strategies.
In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains. Unlike general FedIT, which enhances model effectiveness across diverse tasks without accounting for local data shortages, FedDIT addresses the unique challenge of clients only having limited domain-specific data. To overcome this, FedDIT enriches local data through a specific instruction augmentation strategy, crucial for effective instruction tuning and avoiding performance degradation. Our results reveal that, except for the code domain, where a standardized paradigm is primarily adhered to, relying solely on high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance (see underlined values in Table 2). FedDIT aims to develop a domain-specific LLM through collaborative training and instruction augmentation while safeguarding client privacy, ensuring proficiency in tasks pertinent to its designated domain. For simplification and better instruction quality Zhang et al. (2024b); Toshniwal et al. (2024), we focus on a specific scenario of FedDIT involving a server-hosted public dataset, abstracting open-source instruction datasets on the web across multiple domains. This dataset is used for various instruction augmentation strategies, enhancing model performance in specific domains.
@@
To maximize the cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, Federated Instruction Tuning of LLMs with Domain Coverage Augmentation, termed FedDCA. This algorithm employs a greedy client center selection process and implements instruction augmentation through dense retrieval on the server side. The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain (as shown in Figure 1). By strategically optimizing domain coverage at each step, FedDCA efficiently constructs the augmented train set that enhances both the learning and generalization capabilities of the model, leading to superior performance on domain-specific tasks. Furthermore, to mitigate computational overhead on the client side and enhance the system scalability, we propose FedDCA, which employs heterogeneous encoders of different sizes and capacities. To achieve feature alignment, we train a projector on the server side using public data and employ contrastive learning techniques.
We demonstrate the effectiveness and plug-and-play capability of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical. These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods. Results show that FedDCA outperforms other methods from at least 0.84% to the maximum of 29.19% in terms of model performance, with a relative domain coverage improvement from 4.82% to 21.36%. Extensive experiments across privacy analysis, computational efficiency and other further investigations are also conducted.
@@ 2 Problem Formulation
Federated Domain-specific Instruction Tuning (FedDIT) is a federated learning approach designed to improve the performance of LLMs in specific domains by utilizing limited cross-client private data in combination with domain-specific instruction augmentation strategies Zhang et al. (2024b); Xia et al. (2024); Wang et al. (2022). We abstract the public data as a server-hosted multi-domain dataset (the composition is detailed in Table 1). The presence of a public dataset on the server is only one possible scenario, and the proposed algorithm is independent of the presence of a public dataset on the server. Even if the server does not have a public dataset, clients can retrieve public instructions by themselves based on the received client centers from the server or perform self-instruct, thereby achieving data augmentation.
@@
The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data Ye et al. (2024b); Zhang et al. (2024b, 2023b), which is defined as:
$\small\operatorname*{arg\,min}_{\Delta\phi}\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l},D_{k}^{g};\alpha_{k}\right)\right\},$ (1)
where $F_{k}(\cdot)$ is the accumulated instruction fine-tuning loss of model $P_{\phi+\Delta\phi_{k}}$ computed on the client $c_{k}$’s augmented dataset $D_{k}^{l}\cup D_{k}^{g}$. $p_{k}$ is client $k$’s weight based on data ratio, and $\alpha_{k}=\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$ is the ratio of public data. Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model. The client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ is calculated as:
$F_{k}(\phi,\Delta\phi_{k};\mathcal{D})\triangleq\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}l(P_{\phi+\Delta\phi_{k}};x_{j}),$ (2)
where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$. The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction and $y$ is the corressponding response.
In summary, we formalize FedDIT as a novel FL framework that focuses specifically on domain-specific instruction tuning. In contrast, FedIT typically demands a substantial quantity of instructions for fine-tuning. Domain-specific instruction augmentation can be divided into two categories: 1) Generative methods, which generate instructions through the pre-trained LLM locally or external APIs. 2) Retrieval-based methods, which retrieve instructions from the web. The former methods either compromise privacy or have high computational overhead. Thus, in this work, we focus on retrieval-based methods to utilize diverse and high-quality public data.
@@ 3 What Truly Counts in FedDIT
(a) Exp1. Code
(b) Exp1. Med
(c) Exp1. Fin
(d) Exp1. Math
(e) Exp2. Performance (%)
(f) Exp2. Domain Coverage
Figure 2: (a)-(d) Performance (%) of different heterogeneity in each domain with 10 and 100 clients. (e)-(f) Performance and domain coverage of iid and non-iid settings on different domains. Experiments show that cross-client domain coverage significantly impacts model effectiveness, while data heterogeneity shows no monotonic correlation with performance.
Data Heterogeneity Is Not Matter in FedDIT.
We adopt the Dirichlet distribution Wang et al. (2020); Yurochkin et al. (2019) to construct various heterogeneity ($\alpha=[0.01,0.1,1,10]$) and use k-means with $\xi=100$ to pseudo labeling instructions.
We perform instruction tuning on both 10 and 100 clients with 2 randomly selected clients participating in each round. For each domain, we only use the in-domain data and then perform FedIT. We perform the experiments 3 times with different random seeds (42, 43 and 44) and report the average performance and the standard deviation in each domain. As shown in Figures 2(a), 2(b), 2(c) and 2(d), the performance of LLM does not decrease due to the increase of data heterogeneity but shows a non-monotonic correlation, which indicates that the performance of LLM does not directly depend on data heterogeneity and other factors that play a key role.
Domain Coverage: A Key Factor in FedDIT.
@@
Inspired by the facility location function Cornuejols et al. (1983), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
$\small d(D^{d},D^{c})=\frac{1}{\left|D^{d}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (3)
where $sim(d,v)$ is the cosine similarity between $d$ and $v$. Note that we only use the in-domain data in $D^{c}$ for the domain coverage calculation to avoid misleading.
To better align with the real-world scenarios, we explore instruction augmentation based on both iid and non-iid cross-client data distribution and then perform dense retrieval for FedDIT. We set the number of clients to 10, while each client has 100 local instructions and obtains 1000 augmented public instructions from the server. For the iid setting, we randomly sample 1k from the public dataset and divide them into 10 shards as each client’s local data. For the non-iid distribution, we perform k-means clustering with $\xi=100$. Each client randomly samples 100 instructions from randomly selected distinct clusters.
Figures 2(e) and 2(f) presents the performance and the according domain coverage of FedDIT in different domains with iid and non-iid settings. We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.

### v6

# Optimizing Cross-Client Domain Coverage for Federated Instruction Tuning of Large Language Models

@@ Abstract
Federated domain-specific instruction tuning (FedDIT) for large language models (LLMs) aims to enhance performance in specialized domains using distributed private and limited data, yet identifying key performance drivers and optimal augmentation strategies remains challenging. We empirically establish that cross-client domain coverage, rather than data heterogeneity, is the pivotal factor. We then introduce FedDCA, an algorithm that explicitly maximizes this coverage through diversity-oriented client center selection and retrieval-based augmentation, constructing diverse, non-redundant cross-client instruction sets. Extensive experiments across multiple domains demonstrate FedDCA’s superiority over eleven baselines, achieving performance gains of up to 29.19% and domain coverage improvements of 4.82%-21.36%. FedDCA maintains its effectiveness in diverse and challenging scenarios, including data selection, held-out settings where task-specific public data is scarce and various data heterogeneity, with manageable privacy risks. This work clarifies critical FedDIT dynamics and presents FedDCA as an effective, privacy-preserving, and scalable solution for advancing domain-specific LLM tuning.
Optimizing Cross-Client Domain Coverage for Federated Instruction Tuning of Large Language Models
Zezhou Wang1, Yaxin Du2, Xingjun Ma3, Yugang Jiang3, Zhuzhong Qian1 *, Siheng Chen2 *, 1Nanjing University, 2Shanghai Jiao Tong University, 3Fudan University {zzw.cs@smail, qzz@}nju.edu.cn, {dorothydu, sihengc}@sjtu.edu.cn, {xingjunma, ygj}@fudan.edu.cn
\$*\$\$*\$footnotetext: Corresponding author.
@@ 1 Introduction
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data McMahan et al. (2017); Ye et al. (2024c); Zhang et al. (2023b). This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization. Despite the establishment of various FedIT frameworks Ye et al. (2024c); Kuang et al. (2023); Zhang et al. (2023b), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications. For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts Zhang et al. (2024c).
We explore Federated Domain-specific Instruction Tuning (FedDIT), an innovative approach that harnesses federated learning (FL) to unlock the potential of Large Language Models (LLMs) in specialized domains. A significant hurdle in deploying such models arises when multiple entities hold valuable but limited and privacy-sensitive data—a common scenario in fields like medical diagnosis Guan and Liu (2023); Hu et al. (2023) and financial risk assessment Abadi et al. (2024). These entities require collaborative training of domain-specific LLMs without direct data sharing Xu et al. (2024a); Zhang et al. (2024c). Our findings underscore a key limitation: exclusive reliance on local, in-domain data, despite its quality, often results in subpar performance due to insufficient scale (see Table˜1). FedDIT tackles this by integrating carefully designed instruction augmentation strategies (detailed in Section˜2) that expand local datasets while upholding privacy. This augmentation is not merely a supplement but a crucial enabler for robust instruction tuning and for averting performance decline. Focusing on practical and high-quality augmentation Zhang et al. (2024c); Toshniwal et al. (2024), we investigate a FedDIT setup employing a server-hosted public dataset (a cross-domain instruction set, detailed in Section˜B.1) and utilize various sampling strategies to realize the data augmentation (discussed in Section˜3). This dataset, abstracting diverse open-source instructions, provides a versatile foundation for augmentation techniques designed to elevate model performance within specific target domains.
@@
To maximize cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, FedDCA (Domain Coverage Augmentation). The inability of direct data sharing under privacy constraints (detailed in Section˜5.1) motivates our approach. Initially, each client computes a set of candidate cluster centers from its local data; these serve as privacy-preserving proxies for its semantic distribution. FedDCA then employs a server-side, domain-coverage-oriented selection algorithm (detailed in Section˜5.2) to choose a strategic subset (denoted as client centers) from all uploaded candidate centers. These client centers then guide server-side dense retrieval for instruction augmentation. This selection process serves two critical purposes: 1) it provides a privacy-preserving way to capture the semantic diversity of distributed data, and 2) it enables efficient retrieval of relevant public instructions that complement the collective needs of the clients. By strategically optimizing the cross-client domain coverage, FedDCA efficiently constructs an augmented train set that enhances the model’s generalization capability, leading to superior performance on domain-specific tasks.
To substantiate our claims, we conduct a rigorous empirical study. This includes foundational experiments identifying the key performance drivers in FedDIT, the development and evaluation of our proposed FedDCA algorithm across multiple specialized domains (medical, financial, and mathematical) against eleven baselines, and in-depth analyses of its practical aspects such as privacy, scalability, and robustness in challenging settings (detailed in Sections˜6 and B). Our key contributions are as follows:
  • We empirically reveal a critical finding: in Federated Domain-specific Instruction Tuning (FedDIT), cross-client domain coverage, rather than data heterogeneity, substantially impacts LLM effectiveness. We propose a novel representation-based metric to quantify this cross-client domain coverage.
  • We propose FedDCA, an algorithm that maximizes cross-client domain coverage through diversity-oriented client center selection and retrieval-based instruction augmentation.
  • Extensive experiments across diverse domains demonstrate FedDCA’s superior effectiveness and plug-and-play capability, outperforming baselines by at most 29.19% in model performance and improving relative domain coverage by 4.82% to 21.36%. We also show its privacy-preserving capability in Section˜B.5.
  • We demonstrate FedDCA’s robustness and practical applicability by evaluating its performance in data selection scenarios (achieves comparable performance using only 10% of baseline’s data), held-out settings where the server lack of task-specific public data, and its consistent effectiveness across varying inter-client data heterogeneity (Sections˜6.2, B and C).
2 Related Work
Federated Instruction Tuning.
Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM Zhang et al. (2023c); Wei et al. (2022). Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks Ye et al. (2024c, b); Zhang et al. (2023b) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigated the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
Some progress has been made in the study of FedDIT Zhang et al. (2024c); Wang et al. (2024); Ye et al. (2024a), such as FewFedPIT, which addresses data scarcity by locally generating data using pre-trained LLMs and is the first to explore memory extraction attacks within FedIT. Another relevant approach is FedIT-U2S Ye et al. (2024a), which focuses on enabling FedIT when clients only possess raw, unstructured documents. Similar to our conceptualization of a server-hosted public dataset for augmentation, FedIT-U2S also leverages server-side resources (the example database and the LLM for instruction generation) to enrich client data, albeit with a focus on structuring initially unstructured data rather than augmenting existing instructions with diverse new ones.
Non-IID data distribution is a common challenge in FL Karimireddy et al. (2020); Li et al. (2020); Sattler et al. (2019); Li et al. (2019, 2021). However, the impact of data heterogeneity on FedDIT has not been fully explored. To fully understand the impact of data heterogeneity on FedDIT and what truly matters in FedDIT, we conduct a comprehensive analysis in Section˜4.
Domain Instruction Augmentation.
In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code Nijkamp et al. (2023); Luo et al. (2024), medical Zhang et al. (2023d), financial Yang et al. (2023b, a); Zhang et al. (2023a); Wu et al. (2023), mathematical Yue et al. (2024); Luo et al. (2023)).
Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
Augmentation strategies broadly fall into two categories: those leveraging existing/mined data and those generating new instructions. Methods based on existing or mined data include reusing human-curated public datasets Wang et al. (2023); Zhang et al. (2023e), retrieving from large instruction pools Xia et al. (2024); Jiao et al. (2023), or scaling instruction acquisition from the web Yue et al. (2024); Zhou et al. (2024). These approaches offer access to diverse, potentially high-quality instructions with generally favorable efficiency and privacy, but depend on the availability and relevance of such external data for the target domain. Conversely, generative methods Wang et al. (2022); Zhang et al. (2024c); Wang et al. (2024) create new instructions using LLMs (locally or via APIs). While potentially highly tailored, they often incur significant computational/API costs, raise privacy concerns, and face challenges in ensuring consistent instruction quality and diversity, as detailed in Appendix˜B.
In summary, adapting existing instruction augmentation techniques to federated learning presents significant challenges in balancing data quality, diversity, efficiency, privacy, and scalability. Current methods often involve trade-offs between costly, quality-variable generative approaches and data-dependent retrieval strategies. This necessitates robust augmentation frameworks tailored for FedDIT, motivating our work in this area.
@@ 3 Problem Formulation
Federated Domain-specific Instruction Tuning (FedDIT) is a federated learning approach designed to improve the performance of LLMs in specific domains by utilizing limited cross-client private data in combination with domain-specific instruction augmentation strategies Zhang et al. (2024c); Xia et al. (2024); Wang et al. (2022).
For practical and scalable augmentation within FedDIT, we conceptualize the primary source of augmentation instructions as a server-hosted, multi-domain public dataset (detailed in Section˜B.1). This strategic choice offers several key advantages in a federated context. Firstly, it provides a unified and manageable resource pool, abstracting the complex origins of public instructions (whether curated, mined, or pre-generated). Secondly, server management of this dataset decouples data sourcing from the core federated augmentation logic (such as retrieval strategies), allowing the latter to focus on optimizing instruction selection and distribution. Finally, a centralized public dataset enables efficient server-side pre-processing and indexing, which ensures consistent data quality and supports sophisticated retrieval mechanisms beneficial to all participating clients.
@@
The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data Ye et al. (2024c); Zhang et al. (2024c, 2023b), which is defined as:
$\operatorname*{arg\,min}_{\Delta\phi}\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k};\alpha_{k}\right)\},$ (1)
where $F_{k}(\cdot)$ represents the accumulated instruction fine-tuning loss of model $w_{\phi+\Delta\phi_{k}}$ evaluated on client $c_{k}$’s augmented dataset $D_{k}^{l}\cup D_{k}^{g}$. Here, $p_{k}$ denotes client $k$’s weight based on data ratio, and $\alpha_{k}=\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$ represents the proportion of public data. Equation 1 optimizes the global model by minimizing the weighted sum of empirical losses across clients’ augmented instructions, thereby enhancing in-domain utility. The empirical loss for client $c_{k}$, $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$, is computed as $\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}l(w_{\phi+\Delta\phi_{k}};x_{j})$, where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$ and $l(\cdot)$ is the instruction tuning loss function.
For a client $c_{k}$ with private instruction set $\mathcal{D}_{k}$, we first encode instructions into embeddings using encoder $w_{\mathrm{enc}}$. These $d$-dimensional vectors are then clustered into $\xi$ groups via $k$-means, yielding centroid set $\mathcal{C}_{k}=\{\mathbf{c}_{k,1},\mathbf{c}_{k,2},\ldots,\mathbf{c}_{k,\xi}\}$, which is subsequently transmitted to the server.
We first introduce the Direct Retrieval baseline, which performs independent dense retrieval Zhao et al. (2022) for each client without explicit consideration of cross-client domain coverage. For each centroid $\mathbf{c}_{k,j}$, the server retrieves the top-$\frac{N_{k}^{p}}{\xi}$ most similar public instructions using cosine similarity. The retrieved instructions are aggregated as client $c_{k}$’s augmented data $D_{k}^{p}=\bigcup_{j=1}^{\xi}\mathcal{R}_{k,j}$. Finally, client $c_{k}$ fine-tunes its model $w_{k}$ on the combined dataset $D^{k}=D_{k}^{l}\cup D_{k}^{g}$ of private and retrieved public instructions.
@@ 4 What Truly Counts in FedDIT
(a) Exp1. Med
(b) Exp1. Fin
(c) Exp1. Math
(d) Exp2. Performance (%)
(e) Exp2. Domain Coverage
Figure 1: (a)-(c) Performance (%) of different heterogeneity in each domain with 10 and 100 clients. (d)-(e) Performance and domain coverage of iid and non-iid settings on different domains. Experiments show that cross-client domain coverage significantly impacts model effectiveness, while data heterogeneity shows no monotonic correlation with performance.
4.1 Data Heterogeneity: A Minor Factor
We adopt the Dirichlet distribution Wang et al. (2020); Yurochkin et al. (2019) to construct various heterogeneity ($\beta=[0.01,0.1,1,10]$) and use k-means with $\xi=100$ to pseudo labeling instructions.
We then perform instruction tuning on both 10 and 100 clients with 2 randomly selected clients participating in each round. For each domain, we only use the in-domain data and then perform FedIT. We repeat the experiments for 3 times with different random seeds (42, 43 and 44) and report the average performance and the standard deviation in each domain. As shown in Figures˜1(a), 1(b) and 1(c), the performance of LLM shows a non-monotonic correlation with the data heterogeneity, which indicates that the performance of LLM does not directly depend on data heterogeneity and exist other factors playing a key role.
4.2 Domain Coverage: A Key Factor
@@
Inspired by the submodular function Krause and Golovin (2014), we first define a general coverage metric. For any two sets of item embeddings, $S_{1}$ (the reference set) and $S_{2}$ (the covering set), their coverage, denoted $d(S_{1},S_{2})$, is given by:
$d(S_{1},S_{2})=\frac{1}{|S_{1}|}\sum_{s_{1}\in S_{1}}\max_{s_{2}\in S_{2}}\operatorname{sim}(s_{1},s_{2}),$ (2)
where $\operatorname{sim}(\cdot,\cdot)$ is the cosine similarity between embeddings. Then the domain coverage of $D^{c}$ respect to $D^{d}$ is $d(D^{d},D^{c}\cap D^{d})$. The use of $D^{c}\cap D^{d}$ ensures that we only consider the in-domain portion of the aggregated client data for this particular coverage calculation, preventing misleading scores from out-of-domain client data, consistent with our original formulation.
To better align with our setup, we explore instruction augmentation based on both iid and non-iid cross-client local data distribution and then perform dense retrieval Zhao et al. (2022) for FedDIT (the Direct Retrieval method). We set the number of clients to 10, while each client has 100 local instructions and obtains 1000 augmented public instructions from the server. For the iid setting, we randomly sample 1000 from the public dataset and divide them into 10 shards as each client’s local data. For the non-iid distribution, we perform k-means clustering with $\xi=100$. Each client randomly samples 100 instructions from randomly selected distinct clusters and then performs retrieval for both settings.
Figures˜1(d) and 1(e) presents the performance and the according domain coverage of FedDIT in different domains with iid and non-iid settings. We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.